OpenStudy (anonymous):

Write the ratio of whole numbers using fractional notation. Write the fraction in simplest form. Remember to cancel units when applicable. 66 liters to 102 liters A. 10/11 liters B. 66/60 liters C. 10/11 D. 11/17

OpenStudy (anonymous):

66 liters/102 liters x (1/liters)/(1/liters) = 66/102 x (1/6) / 1/6) = 11/17 So "D" is the answer

OpenStudy (anonymous):

Not sure the logic is sound with the guy above. But 66/102 is the fraction. Divide by a form of 1 equal to 2/2 and you get 33/51. Both are divisible by 3 so we divide again by a form of one 3/3. 33/3=11 and 51/3=17. So the simplest fraction is 11/17 which is D.

OpenStudy (anonymous):

Thanks